Grade 10General Physics

A hollow sphere of external and internal diameter 4cm and 2 cm , respectively floats ina liquid of density 3.5g/cm^3. The level of the liquid coincides with the centre of the sphere . Calculate the density of the material of the sphere

Bhavan
8 Years agoGrade 10

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear Bhavan
Volume of sphere = (4/3) * pi* (r1³ - r2³)
= (4/3) *pi* (4³ -2³) = 234.57 cm³
Now according to the question half of the sphere is floating and other half of the sphere is immersed in the liquid.
So , weight of the floating body = weight of the liquid displaced
Volume of sphere floating * density of sphere * g = volume of sphere immersed * density of liquid * g
234.57/2 * ds * g = 234.57/2 * 3.5 * g
ds = 3.5 gm/cm³
Regards
Arun (askIITians forum expert)